OpenStudy (anonymous):

okay, so i have to change the equatiion into slope intercept form and the equation is 3(x-2)=5(y+4).... I got y=3/5x - 26/5.... does that look right.... Note: Slope intercept form is y=mx+b okay, so i have to change the equatiion into slope intercept form and the equation is 3(x-2)=5(y+4).... I got y=3/5x - 26/5.... does that look right.... Note: Slope intercept form is y=mx+b @Mathematics

OpenStudy (anonymous):

3x - 6 = 5y + 20 3x - 26 = 5y 3/5x - 26/5 = y yes you are correct

OpenStudy (anonymous):

YAY! thank you so much!